# KinesCeTI: A Modular and Size-Adaptable Force Feedback Glove with Interchangeable Actuation for the Index and Thumb

### v1

@@ III-C Linkage system
In both fingers, an assembly consisting of a Knuckle Hub, housing both actuators and sensors and the tendon-linkage system, is mounted on the abduction/adduction shaft. The assemblies are almost identical, differing only in linkage lengths, whose calculation is explained in this subsection. Correct sizing is necessary in order to enable adaptation to different hand sizes. Numerous publications describe the variability of human hand motion and dimensions, using in this work various studies as references [chen2011human, mclain2010use, ueki2010development, garrett1971adult, ozsoyattempt, ash1996proximal, kong2017investigation, jones2006human]. However, it should be noted that these datasets do not represent the entire human population, as they are limited to specific countries, age groups and professions.
The design objective is therefore to create a structure that adapts to most human hands across their full range of motion. To this end, a 2D geometrical analysis is conducted, where both fingers and linkages are modeled as lines without thickness. The chosen structure consists of two linkages in series with revolute joints, with the first joint at the Knuckle Hub and the last at the thimble. Regarding their geometries, the first linkage is a straight beam, while the second is optionally arched. Input parameters include three phalanx sizes (small, medium, large), finger joint angles (tested in 10° steps) and thimble attachment either at the fingertip or the second phalanx, with 20 mm offset perpendicular to their middle points. Values are shown in Tables I and II. Linkage combinations are tested for lengths of 60–200 mm and 10 mm steps for both links. Each combination is evaluated for coverage of the motion range and potential collisions with the finger. When these occur, an arched second linkage (up to 30 mm height) is tested. Combinations failing both checks are rejected. The overall testing process is shown in Fig. 3.
TABLE I: Finger Joint Motion Ranges and Start Joint Coordinates
Finger MCP/CMC PIP/MCP DIP/IP Start joint
Index $90^{\circ} 0$$110^{\circ}$ $90^{\circ} [0$, 45][0,\,45] $m$
Thumb $70^{\circ} 0$$80^{\circ}$ $90^{\circ} [0$, 70][0,\,70] $m$
TABLE II: Phalanx Length Sizes ($m$)
@@
Figure 3: Linkage algorithm: (a) successful configuration with variables indicated, (b) invalid configuration due to insufficient linkage length, (c) invalid configuration caused by intersection with the finger at fingertip attachment and (d) valid configuration using a modified arched geometry for the second linkage.
Figure 4: Algorithm results for (a) index and (b) thumb. Green tiles mark successful combinations, with values indicating the arched height.
Among feasible solutions (Fig. 4), the shortest configurations are selected to minimize weight and volume: L1 = 170 mm, L2 = 130 mm and 30 mm height for the index and L1 = 160 mm, L2 = 90 mm and 20 mm height for the thumb. To preliminarily evaluate force capabilities, load tests with perpendicular forces were applied to the linkages. For the first linkage, only the index configuration (L1 = 170 mm, corresponding to the longest case) was tested, withstanding a maximum load of 52.6 N. For the second linkage, the index configuration (L2 = 130 mm) withstood 45.3 N, while the thumb configuration (L2 = 90 mm) sustained 62.5 N. These tests were performed without additional structural components (bolts, pulleys, or tendons) and represent preliminary feasibility checks rather than definitive strength measurements.
@@ III-D Modular thimbles
The second linkage terminates in the thimble assembly, which serves as the end effector and provides two revolute joints: one for flexion/extension and one for abduction/adduction. Although the latter is already present at the knuckle joint, this extra DOF is required to maintain alignment with the finger’s axis, compensating for the plane misalignment between finger motion and linkage orientation. The thimble assembly includes an attachment system consisting of a slider–lock mechanism, with an insertion slot and a compliant PETG latch that secures it in place. This system is shown in Fig. 5(a) and its operation in Fig. 5(b).
Two types of thimbles were developed. The first is designed for fingertip attachment, which requires customized shapes because the fingertip is a free end lacking a joint to secure the thimble and exhibits greater variability in geometry (Fig. 5(c). For flexibility, the chosen material is TPU 95A. These thimbles are produced in ten sizes for each finger with inner diameters from 14 to 23 mm (1 mm step). Thumb thimbles include additional width variations (2-6.5 mm extra), resulting in a total of 20 designs. All designs are illustrated in Fig. 5(e). The second type is a single-size module for the second phalanx, consisting of a polypropylene strap secured by a PETG hub with an integrated spring latch, which releases when pressed (Fig. 5(d)).
To evaluate the force capabilities of the thimble assembly, preliminary load tests were conducted on both types. For the strap design, five samples were tested, withstanding loads of $98.98\pm 14.55$ N before failure, which consisted of strap loosening while remaining operational. For the TPU fingertip thimbles ten samples were tested, withstanding $45.42\pm 6.11$ N. In this case, failure occurred when the TPU slider pulled out of the PETG holder, after which the thimble remained functional.
Figure 5: Thimble system: (a) compliant latch design and placement, (b) latch operation, (c) TPU fingertip thimble and attachment, (d) strap thimble and attachment and (e) TPU fingertip thimble designs for index and thumb.
@@ III-E Force transmission system
The linkages, thimbles and base structure form a continuous mechanical chain connecting the fingers to the dorsal side of the hand. However, this system is mechanically passive and therefore requires additional components for sensing and force feedback. To minimize weight and inertia on the fingers, both actuation and sensing elements are placed at the Knuckle Hubs. Motion and force are transmitted through a bidirectional closed-loop tendon-pulley architecture, also known as a cable-routed pulley (CPR) configuration with endless cables [grosu2018driving]. In the current design, two tendons made of braided polyethylene line (Spiderwire Dura-4 Braid, 0.35 mm diameter, 35 Kg rated load) are used. Each tendon is anchored at two points: the ratchet-pawl pretensioner on the second linkage, where pretensioning occurs and the main pulley at the Knuckle Hub. The tendon is secured at both ends with constrictor knots, which are reinforced with multiple loops to ensure reliability.
Figure 6: Side view of the index linkage showing the force transmission system and enabled DOFs.
The tendon routing is designed to ensure that the close loop system transmits the angle of both linkages to the main pulley, while intermediate pulleys guide the tendons within the structure and prevent slack, as illustrated in Fig. 6. For sensing, two AS5600 encoders are used: one at the first joint to measure the angle of the first linkage (L1) and another at the main pulley to capture the combined angle of both linkages, thus providing two measured DOFs.
@@ III-F Actuation system
III-F1 Ratchet pawl mechanism
The first actuation system is a custom compliant ratchet pawl-mechanism (Fig. 7(b)). The ratchet is fixed to the main pulley, while the pawl with a built-in spring is attached to the servomotor. This mechanism provides binary feedback: the pulley is either free to rotate or blocked by the pawl when actuated by the servomotor. It thus enables unidirectional stiff boundary force feedback with high backdrivability. Stiffness is inherently high due to its ratchet-pawl design, although slightly reduced by the compliance of the mechanical elements in series in the actuation chain.
A trade-off exists between resolution, robustness and speed: more teeth increase angular resolution and clutching speed, but reduce robustness. 3D printing constraints also limit possible geometries. In our design, the ratchet had a maximum outer diameter of 16 mm and an inner diameter of 12–13 mm (tooth thickness 1.5-2 mm). Among the tested options (12–30 teeth), the 20-tooth configuration ($18^{\circ}$ resolution/tooth) was selected as the best balance of resolution and strength, with preliminary load tests on three samples yielding relatively high failure loads of 155.04 N, 157.1 N and 131.8 N, consistently above the maximum thimble loads.
Figure 7: Modular actuation system: (a) main pulley socket, (b) ratchet–pawl mechanism, (c) main pulley without feedback, (d) main pulley with ratchet–pawl mechanism and (e) main pulley with one-way clutch.
The mechanism is lightweight (pawl, ratchet and screw weigh 1.74 g) and simple. In addition to the Knuckle Hub design, modified versions were integrated into the thumb structure to provide abduction–adduction feedback (Fig. 2(a)-(b)). Clutching latency was characterized for each subsystem, with 20 samples collected per motor. For the Knuckle Hubs, it yielded a mean latency of $64.475\pm 22.99$ ms (range: 9–108 ms) for the combined 40 samples. For the thumb base, the fixed servomotor showed a latency of $58.5\pm 27.98$ ms (range: 9–108 ms), while the bevel servomotor exhibited $74.25\pm 28.96$ ms (range: 26–121 ms).
@@ III-F2 One-way clutch mechanism
To enable active unidirectional feedback with variable impedance while maintaining good backdrivability, a novel clutch mechanism is developed. In this design, both clutching and feedback are driven by the same servomotor, using an architecture similar to [romeo2021passively] but based on a compliant ratchet–pawl principle. The components, shown in Fig. 8(a), are as follows:
@@
  • Pawl Carrier (PC): surrounds the ORS and houses a compliant pawl.
  • Driving Drum (DD): encloses both the PC and ORS and it’s concentric to them via a bearing and a screw. In this design, it is driven by the servomotor through a direct drive gear. Its inner profile has an irregular geometry that engages the pawl.
@@
Figure 8: One-way clutch mechanism: (a) frontal and exploded view of the assembly, (b) clutching operation.
@@
  • Clutching: when feedback is required, the motor drives the DD while the FR holds the PC. The inner profile of the DD has a decreasing radius in the pawl area, which pushes the pawl inward to engage the ratchet.
  • Feedback: once engaged, the system enables active feedback, capable of applying resistance or active forces against the finger.
  • Unclutching: reverse motor rotation or opposite finger motion disengages the pawl, restoring free motion.
@@
The clutch mechanism was implemented in the exoskeleton and measured in both the index and thumb Knuckle Hubs, with 20 samples collected in each. The exoskeleton was kept still during the measurements and was not attached to any hand, to avoid finger movements that may affect the latency. Times were measured between the sent command and motion detection using the corresponding AS5600 sensors. For the combined data, mean latency was $88.7\pm 24.68$ ms (range: 17–152 ms) and mean clutching angles were $52.27\pm 14.57^{\circ}$ (range: 19.46–87.12). The complete clutch module, including bearing and bolt, weighs approximately 4.39 g and measures $\diameter 18.6\times 14$ mm (excluding the servomotor).
@@ III-G Control system
The control system uses an ESP32 in the Main Hub to drive the servomotors and acquire sensor data. AS5600 encoders provide analog PWM outputs that are read through voltage dividers, while servomotor potentiometer signals are buffered with voltage followers to ensure stable readings. Sensing and control electronics are powered via USB from the host computer, whereas the four servomotors are powered separately at 6 VDC (up to 1 A) using a Voltcraft SNG-1000-OC. During experiments, the ESP32 handled sensing and motor control while communicating over serial with a CHAI3D simulation [conti2003chai] running on the host computer, which computes interactions with virtual objects. Further details are presented in the next section.
@@ III-H Overall summary
This section presented the design of a modular and adaptive exoskeleton weighing approximately 376 g, including TPU thimbles and servomotors but excluding the previously described actuation mechanisms. The system integrates interchangeable actuation modules, modular thimbles and an adaptive linkage structure to fit different hand sizes. Its sensing and actuation architecture enables both binary and variable force feedback. These features are evaluated in the user studies described in the next section.

### v2

@@ III-C Linkage system
For both digits, a Knuckle Hub assembly housing the actuators, sensors, and tendon–linkage system is mounted on the ab-/adduction shaft. The two assemblies are mechanically identical, differing only in linkage lengths, whose calculation is explained in this subsection. Correct sizing is required to enable adaptation to different hand sizes. Human hand motion and dimensional variability are well documented in the literature, and multiple references are used [9, 17, 35, 16, 29, 5, 26, 25]. It should be noted, however, that such datasets do not cover the entire human population, as they are limited to specific countries, age groups, or professions.
The objective is to identify linkage length combinations that accommodate most human hands across the full range of digit motion and allow attachment either at the fingertip or the second phalanx, enabling a general-purpose glove. The mechanism consists of two serial linkages connected by revolute joints: the base joint at the knuckle hub (ExobasejointExo\ base\ joint), an intermediate linkage connection (CommonjointCommon\ joint), and a distal joint at the thimble attachment point (ThimblejointThimble\ joint). The first linkage ($L1$) is modeled as a straight beam, while the second linkage ($L2$) may be straight or arched.
The analysis considers three representative phalanx size sets, spanning the minimum, mean, and maximum finger dimensions reported in the literature and denoted as smallsmall, mediummedium, and largelarge (Table I). Digit joint angles are sampled in $10^{\circ}$ increments over their physiological ranges (Table II). The thimble is assumed to be attached at either the fingertip or the second phalanx, with a fixed perpendicular offset of 20 mm (thimble height). Candidate linkage length combinations $(L1,L2)$ are evaluated over 60–200 mm in 10 mm steps. For each linkage length combination and digit configuration, the algorithm first verifies whether both linkages can geometrically form a common joint, and then checks whether the resulting configuration spans the full joint angle range without intersecting the digit geometry. If a collision is detected (Fig. 4(c)), an alternative configuration with an arched second linkage is evaluated, with arch heights up to 30 mm in 10 mm steps. Linkage combinations that fail to form a valid joint ((Fig. 4(b))) or to satisfy the motion coverage and collision constraints across all configurations are rejected.
TABLE I: Phalanx Length Sizes ($m$)
@@
TABLE II: Digit Joint Motion Ranges and Base Coordinates
Finger MCP/CMC PIP/MCP DIP/IP Base joint
Index $90^{\circ} 0$$110^{\circ}$ $90^{\circ} [0$, 45][0,\,45] $m$
Thumb $70^{\circ} 0$$80^{\circ}$ $90^{\circ} [0$, 70][0,\,70] $m$
Figure 4: Linkage algorithm: (a) successful configuration with variables indicated, (b) invalid configuration due to insufficient linkage length, (c) invalid configuration caused by intersection with the digit at tip attachment and (d) valid configuration using a modified arched geometry for the second linkage.
Figure 5: Algorithm results for (a) index and (b) thumb. Green tiles denote successful combinations and their arc height ($m$).
From feasible solutions (Fig. 5), configurations with the shortest combined length were chosen to minimize weight and volume: L1 = 170 mm, L2 = 130 mm and 30 mm height for the index, and L1 = 160 mm, L2 = 90 mm and 20 mm height for the thumb. To preliminarily assess force capability, quasi-static load tests were conducted using a custom test bench. Each linkage was rigidly clamped at one end, while a perpendicular tensile load was incrementally applied at the distal end via a Spiderwire Dura-4 Braid cable (0.35 mm diameter, 35 kg rated load), connected to a mechanical tuning-gear mechanism and measured using a calibrated scale gauge. The maximum sustained force prior to structural failure was recorded.
For the first linkage, only the index configuration (L1 = 170 mm, corresponding to the longest case) was tested, withstanding an ultimate tensile load of 52.6 N. For the second linkage, the index configuration (L2 = 130 mm) withstood 45.3 N, while the thumb configuration (L2 = 90 mm) sustained 62.5 N. These tests represent preliminary feasibility checks rather than definitive strength measurements.
@@ III-D Modular thimbles
The second linkage terminates in a thimble assembly that serves as the end effector and provides two revolute joints for flexion/extension and ab-/adduction. Although the latter is already present at the knuckle joint, this additional DOF at the thimble ensures alignment with the digit axis, compensating for plane misalignment between digit motion and linkage orientation. The thimble assembly incorporates a slider–lock attachment mechanism comprising an insertion slot and a compliant PETG latch that secures it in place (Fig. 6(a)–(b)).
Two thimble designs were developed. The first is intended for fingertip attachment and requires customized geometries, as the fingertip is a free end without an anatomical joint for fixation and exhibits higher geometric variability (Fig. 6(c)). For flexibility, the chosen material is TPU 95A, being produced in ten sizes for each digit with inner diameters from 14 to 23 mm (1 mm step). Additional width variations (2–6.5 mm) are implemented for the thumb (Fig. 6(e)). The second design is a single-size module for the second phalanx, consisting of a polypropylene strap secured by a PETG hub with an integrated spring latch, which releases when pressed (Fig. 6(d)).
To evaluate the force capabilities of the thimble assembly, preliminary load tests were performed on both designs using the same test bench described in the previous subsection. For the strap-based design, five samples were tested, withstanding loads of $98.98\pm 14.55$ N before failure due to strap loosening while remaining operational. For the fingertip thimbles ten samples were tested, sustaining $45.42\pm 6.11$ N. In this case, failure occurred when the TPU slider disengaged from the PETG holder, after which the thimble remained functional.
Figure 6: Thimble system: (a) compliant latch design and placement, (b) latch operation, (c) TPU fingertip thimble and attachment, (d) strap thimble and attachment, and (e) TPU fingertip thimble designs for index and thumb.
@@ III-E Force transmission system
The linkages, thimbles and base structure form a continuous mechanical chain connecting the digits to the dorsal side of the hand. As this structure is mechanically passive, additional components are required for sensing and force feedback. To minimize weight and inertia at the digits, both actuation and sensing elements are placed at the Knuckle Hubs. Motion and force are transmitted through a bidirectional closed-loop tendon-pulley architecture, also known as a cable-routed pulley (CPR) configuration with endless cables [18]. This architectural choice enables the implementation of the feedback dimensions from Section 2 (passive/active, uni-/bidirectional, variable impedance). Two braided polyethylene tendons (Spiderwire Dura-4 Braid) are used, each anchored at the main pulley on the Knuckle Hub and at a ratchet–pawl pretensioner on the second linkage, where pretensioning is applied.
The tendon routing transmits the rotation of both linkages to the main pulley, while intermediate pulleys guide the tendons within the structure and prevent slack, as illustrated in Fig. 2. In the absence of force feedback, linkages $L1$ and $L2$ are kinematically independent and can move freely. When force feedback is engaged, the tendon–pulley system kinematically couples their flexion/extension motion. For sensing, two AS5600 encoders are used: one at the first exoskeleton joint to measure the angle of the first linkage ($L1$), and one at the main pulley to capture the combined angle of both linkages, providing two sensed DOFs.
@@ III-F Actuation system
While the described subsystems allow all force feedback dimensions, the actuation system ultimately determines which can be realized. To this end, the Knuckle Hub features a modular actuation architecture that supports motor systems with varying capabilities. The main pulley includes a socket for attaching external actuators or sensing modules (Fig. 7(a)). In this work, two actuation systems were developed and implemented: a ratchet-pawl mechanism and a novel one-way clutch (Fig. 7(c)–(e)), both driven by servomotors, specifically the Bluebird BMS-101DMG, selected based on prior acoustic annoyance studies for wearable actuators [3, 1].
@@
III-F1 Ratchet-pawl mechanism
The first actuation system is a custom compliant ratchet-pawl mechanism (Fig. 7(b)). The ratchet is fixed to the main pulley, while the pawl with a built-in spring is attached to the servomotor, providing binary feedback, allowing the pulley to rotate freely or be blocked. This enables unidirectional stiff boundary force feedback with high backdrivability. Stiffness is inherently high due to its ratchet-pawl principle, but slightly reduced by the overall structure compliance. A trade-off exists between angular resolution, robustness, and clutching speed, as increasing the number of teeth improves resolution but reduces structural strength and is constrained by additive manufacturing limits. With an outer diameter of 16 mm and an inner diameter of 12–13 mm (radial tooth height 1.5–2 mm, face width 4 mm), a 20-tooth configuration ($18^{\circ}$/tooth) was selected from tested designs (12–30 teeth) as the best compromise. Preliminary load tests on three samples yielded failure loads of 155 N, 157.1 N and 131.8 N, consistently above the maximum thimble loads.
Figure 7: Modular actuation system: (a) main pulley socket, (b) ratchet–pawl mechanism, (c) main pulley without mechanism, (d) installed ratchet–pawl, and (e) installed one-way clutch.
The mechanism is lightweight (pawl, ratchet, and screw: 1.74 g) and simple, also adapted for thumb ab-/adduction feedback (Fig. 3(a)-(b)). Clutching latency was characterized with 20 samples per motor, yielding means of $83.15\pm 33.13$ ms (28–150 ms) for the combined 40 Knuckle Hub samples, $71.65\pm 29.07$ ms (37–144 ms) for the fixed thumb motor, and $79.15\pm 20.28$ ms (50–125 ms) for the bevel motor.
@@ III-F2 One-way clutch mechanism
To enable active unidirectional feedback with variable impedance while maintaining good backdrivability, a novel clutch mechanism is developed. In this design, both clutching and feedback are driven by the same servomotor, using an architecture similar to [2] but based on a compliant ratchet–pawl principle. The components, shown in Fig. 8(a), are as follows:
@@
  • Pawl Carrier (PC): surrounds the ORS and houses a compliant pawl with an increasing radial profile.
  • Driving Drum (DD): encloses both the PC and ORS and is concentrically mounted via a bearing and a screw. It is driven by the servomotor through a direct drive gear.
@@
Figure 8: One-way clutch mechanism: (a) frontal and exploded views, (b) Free motion status, (c) Clutching operation, and (d) Feedback mode.
@@
  • Clutching: when feedback is required, the motor drives the DD while the FR holds the PC. During relative rotation, the fixed inner radius of the DD interacts with the pawl’s increasing radial profile, inducing a radial inward displacement that engages the ratchet.
  • Feedback: once engaged, the system enables active feedback, capable of applying resistance or active forces against the digit.
  • Unclutching: reverse motor rotation or opposite digit motion disengages the pawl, restoring free motion.
@@
The clutch mechanism was implemented in the exoskeleton and measured in both the index and thumb Knuckle Hubs, with 20 samples collected in each. The exoskeleton was kept still during the measurements and was not attached to any hand, to avoid finger movements that may affect the latency. Times were measured between the sent command and motion detection using the corresponding AS5600 sensors. For the combined data, mean latency was $100.1\pm 27.78$ ms (range: 27–160 ms) and mean clutching angles were $62.66\pm 15.2^{\circ}$ (range: 29.46–97.44). The complete clutch module, including bearing and bolt, weighs approximately 4.39 g and measures $\diameter 18.6\times 14$ mm (excluding the servomotor).
TABLE III: Hand anthropometric measurements
ID Description Mean SD Min Max ID Description Mean SD Min Max
1 I-DP length 25.04 1.81 21.3 28.58 15 I-DP circumference 50.5 3.51 44 57
2 I-MP length 25.24 2.80 20.98 32.53 16 I-DIP circumference 54.8 2.80 49 60
3 I-PP length 43.66 3.84 36.47 51.61 17 I-MP circumference 59.05 2.80 54 64
4 I-MC length 73.38 3.69 66.37 79.66 18 I-PIP circumference 66.5 2.64 62 71
5 I PIP-crease length 25.61 2.35 20.29 30.15 19 I-PP circumference 69.5 3.15 62 74
6 MI-MC length 67.11 4.66 59.1 75.38 20 T-DP circumference 61.45 3.85 55 69
7 RI-MC length 63.07 3.35 57.59 69.13 21 T-IP circumference 67.25 3.52 61 73
8 LI-MC length 63.05 5.43 54.73 76.62 22 T-PP circumference 69.6 3.29 62 75
9 T-DP length 31.36 2.44 26.6 35.45 23 Knuckles perimeter 205.1 11.09 174 221
10 T-PP length 34.91 4.29 28.08 42.3 24 Palm perimeter 218.65 12.30 189 233
11 T-MC length 50.23 4.50 42.81 59.54 25 Wrist perimeter 174.25 8.73 159 188
12 Hand breadth (MCP) 83.89 3.94 74.32 91.15 26 I length to knuckle* 93.93 7.32 80.31 107.61
13 Hand breadth (palm) 86.52 4.06 76.97 91.87 27 I length to crease* 75.88 5.69 64.34 82.79
14 Wrist breadth 60.25 3.77 52.01 69.72 28 T length to wrist* 116.5 8.06 101.88 130.07
Units: mm. Abbreviations: I=index, T=thumb, MI=middle, RI=ring, LI=little, DP=distal phalanx, MP=middle phalanx, PP=proximal phalanx, MC=metacarpal, MCP/PIP/DIP/IP=standard joints. *Derived segments, calculated as in Fig. 9.
Figure 9: Dimensions of the hand, according to Table III.
@@ III-G Control system
The control system is implemented on an ESP32 DEVKITV1 embedded in the Main Hub, which drives the servomotors and acquires sensor data. Position sensing is provided by seven AS5600 encoders and buffered servomotor potentiometers, while hand orientation is tracked using a 9-DOF IMU (Adafruit BNO055). The system does not include force sensors, and all control algorithms rely exclusively on position measurements from the encoders and servomotor potentiometers. Electronics are powered via USB from the host computer, whereas the four servomotors are powered separately at 6 VDC (up to 1 A) using a Voltcraft SNG-1000-OC. During experiments, the ESP32 communicates over serial with a CHAI3D simulation [13] running on the host computer, which performs digit calibration, computes interactions with virtual objects, and provides position-based feedback commands. Further details are presented in the next section.
@@ III-H Overall summary
This section presented the design of a modular and adaptive exoskeleton with a total mass of 376 g, measured with TPU thimbles and servomotors but excluding actuation mechanisms. The system integrates interchangeable actuation modules, modular thimbles and an adaptive linkage structure to fit different hand sizes. Its sensing and actuation architecture enables both binary and variable force feedback, supporting uni- and bidirectional interaction. These features are evaluated in the user studies described in the next section. To facilitate reproducibility, the CAD models, bill of materials, and assembly documentation of the proposed system are publicly available at https://doi.org/10.5281/zenodo.19219658.